Use the sample data and confidence level given below to complete parts (a) through (d). A research institute poll asked respo1.8 Positive z Scores Standard Normal (2) Distribution: Cumulative Area from the Left 0 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.
Answer #1

Solution :

Given that,

a)

Point estimate = sample proportion = \hat p = x / n = 540 / 924 = 0.584

1 - \hat p = 1 - 0.584 = 0.416

Z\alpha/2 = 1.645

b)

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.645 * (\sqrt((0.584 * 0.416) / 924)

Margin of error = E= 0.027

c)

A 90% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.584 - 0.027 < p < 0.584 + 0.027

0.557 < p < 0.611

The 95% confidence interval for the population proportion p is : , 0.611

d)

option d) is correct